OpenStudy (anonymous):

Simplify and show work please! ^3sqrt. 4x^3y^5 times ^3sqrt. 12y^2 Thanks Simplify and show work please! ^3sqrt. 4x^3y^5 times ^3sqrt. 12y^2 Thanks @Mathematics

OpenStudy (anonymous):

Can you please explain your notation?

OpenStudy (anonymous):

\[\sqrt[3]{4x^3y^5}\times \sqrt[3]{12y^2}\]

OpenStudy (anonymous):

Yayyy! [:

OpenStudy (anonymous):

idk if this steps are right. ^3sqrt. x^3y^3 times ^3sqrt. 4y^2 times ^3 sqrt. 12y^2

OpenStudy (anonymous):

since you have to isolate for x or y, you decide which first and use the one you didnt isolate inside of the equation to later evaluate. Since you wrote all that, I dont quite get the steps you did....

OpenStudy (anonymous):

okay. I hope the server is back.\[\sqrt[3]{x^3y^3y^3}\times \sqrt[3]{4y}\times \sqrt[3]{12}\]

OpenStudy (anonymous):

I know the answer is\[2xy^2\sqrt[3]{6y}\] but how do you get that?

OpenStudy (anonymous):

How did you do it and do not know how you got there? lol

OpenStudy (anonymous):

I separated all the ^3 from the numbers/variables that couldnt be ^3. So\[\sqrt[3]{x^3y^3y^3}\] are separated from\[\sqrt[3]{4y} \times \sqrt[3]{12}\] and so from that info the answer is :\[2xy^2\sqrt[3]{6y}\] does it make sense or no?